OpenStudy (anonymous):

Factor these trinomials with a POSITIVE CONSTANT ( + C) x^2+ 25 a. (x + 5)(x – 5) b. (x + 5)(x + 5) c. (x – 5) (x – 5) d. cannot be factored. a, (x + 5)(x – 5)???? answer.

OpenStudy (anonymous):

d is correct.

OpenStudy (vishal_kothari):

d

OpenStudy (anonymous):

did u get y the answer is d prysik?

OpenStudy (anonymous):

If you are unsure you can always expand the answers a. (x + 5)(x – 5) = x^2-25 b. (x + 5)(x + 5) = x^2+10x+25 c. (x – 5) (x – 5) = x^2-10x+25 Since none of these equal x^2+25 the answer must be d.

OpenStudy (anonymous):

good explaination Zed

OpenStudy (anonymous):

thanks